# Reduced Order Modelling, Dynamical Systems, Deep Learning, Uncertainty Quantification journal: Journal name

### v2

@@ 4 Uncertainty quantification
Most conventional reduction strategies rely on deterministic estimation, where the latent space projection is represented by a single point estimate. This approach fails to capture the inherent uncertainty within the model. In contrast, UP-dROM utilizes variational methods to assess uncertainties, providing a probabilistic projection in the latent space instead of a deterministic one. This is achieved through the use of a Variational Autoencoder (VAE). Specifically, given the observed data $X=D$ (i.e., the data set $D$), the method seeks to infer the posterior distribution $p(z|X=D)$, where $z$ represents the latent variables.
The central idea connecting variational inference and uncertainty quantification revolves around approximating the true posterior distribution $p(z|X=D)$ using a surrogate distribution $q(z|X=D)$. In this case, the surrogate is chosen to be a Gaussian distribution due to its desirable mathematical properties and ease of interpretation. To ensure that the surrogate closely approximates the true posterior, we minimize the Kullback-Leibler Divergence (KLD) between $p(z|X=D)$ and $q(z|X=D)\sim\mathcal{N}(\mu,\sigma^{2})$, as:
@@
Since $p(z|X=D)$ is not known a priori, we instead utilize the Evidence Lower Bound (ELBO), which can be computed at any stage. Variational inference thoery states that minimizing the ELBO, also minimizes the KLD between the true posterior and the surrogate distribution [31]. The ELBO is given by:
@@
where $p(z)$ is the prior Gaussian distribution representing the distribution of the latent space. It is typically chosen as $p(z)\sim\mathcal{N}(0,I)$, which imposes a standard normal prior on the latent variables. This choice ensures that the latent space is continuous and well structured. It also simplifies the mathematical treatment of the model, making the optimisation more tractable. The KLD between two Gaussian distributions such as $q(z|X=D)$ and $p(z)$ can be computed as follows,
@@ 4.1 Uncertainty in space
We can use the uncertainty measure $\nu_{d,t,\xi}$ to visualize how uncertainty is distributed across the computational domain at different time steps for a fixed parametrisation. Figure 9-a shows the $u$-velocity fields at three key time steps: (1) before the wake forms, during the unstable stationary-like behaviour, (2) during the transition, and (3) once the system has reached the stable limit-cycle, as highlighted in the kinetic energy plot. Figure 9-b illustrates the distribution of uncertainty in space $\nu_{d}$ at these same time steps, computed as:
@@
Figure 9: U velocity field at the selected snapshots (a), corresponding uncertainty fields $\nu_{d}$ (b), and the snapshot locations indicated on the kinetic energy signal (c), for a fixed parametrisation $\xi$: $R_{e}=90$.
This figure illustrates a strong correlation between the known space-time dynamics and the uncertainty measurement. Prior to the system’s transitioning to the stable limit cycle, uncertainty is evenly distributed along the symmetry axis of the flow domain, as expected, but this distribution breaks down as the dynamics transition. It is in this transient regime, that the uncertainty across the entire domain is the highest. The highest confidence is observed upstream of the obstacle, where the dynamics remain relatively stationary throughout the time window. In contrast, lower confidence is concentrated at the edges of the obstacle’s wake, suggesting that uncertainty is more closely related to fluctuations in the flow, rather than the instantaneous flow values in specific regions. These fluctuations increase variability within the system, leading to higher uncertainty.
@@ 4.2 Uncertainty in Time
Notably, the true signal lies almost entirely within the confidence interval. It is important to emphasize that this Uncertainty Quantification (UQ) strategy is not time-dependent. If it were, the uncertainty would gradually increase as the dynamic process evolves autoregressively. Instead, we propose a likelihood measure for obtaining a given value, independent of previous values. This approach provides an instantaneous uncertainty measure with minimal computational cost.
As shown in Figure 10, the greatest uncertainties occur at the peaks of the signal, corresponding to instances when the local dynamics are most critical. This typically happens when the obstacle’s wake shifts within the flow domain, causing substantial changes that the model must predict with accuracy. These transitions trigger increased uncertainty. This behavior has a significant physical interpretation: uncertainty is highest in regions and time periods where the flow dynamics are most sensitive to changes, such as during sharp transitions, shifts in wake structure, small-scale patterns, and rare events.
@@
Furthermore, as expected, higher uncertainty is observed in the transient part of the signal compared to the limit cycle. This phenomenon is formally illustrated in Figure 11. To analyse temporal variations in uncertainty for a fixed parameter, we compute the spatially averaged uncertainty $\nu_{t}$ at each time step $t$ over $n_{xy}$ spatial locations,
@@
The width of the bell curve in Figure 11 aligns with the transient phase of the dynamic evolution visible on the kinetic energy signal. Starting from the initial condition, the uncertainty increases as the system transitions towards the limit cycle. Once this stable state is reached, the uncertainty decreases and stabilizes at a plateau. This can be explained by two factors: (1) the regime transition induces variability, which like any other source of variability, is associated with lower confidence, and (2) the training distribution contains fewer transient states compared to the limit cycle. As a result, the model has been trained more extensively on the stable state than on the transient phases.
@@ 4.3 Uncertainty in the Parameter space
Uncertainty Quantification also provides a confidence measure across the parameter space, aggregating the spatial and temporal component,
@@
with $n_{t}$ the time discretization and $\bar{\phi}_{\xi,d,t}$ the mean of the ensemble of size $n$.
As shown in Figure 7, the model can learn and infer within the 2D parameter space. Specifically, it can interpolate within the training subspace and extrapolate with varying accuracy depending on the region of the parameter space. The further the inference distribution deviates from the training distribution, the greater the uncertainty of the model, although this is not uniform. The direction in which the inference distribution moves away from the training distribution affects the performance of the model. Our uncertainty quantification method can measure this non-Euclidean distance. Figure 12 illustrates how our uncertainty quantification process allows us to identify these regions during inference, thus reducing the need for simulation data for evaluation. To improve the performance of the model, it could be retrained using samples from and only from these uncertain regions. An important observation is the strong agreement between the a-posteriori performance evaluation shown in Figure 7 and the a-posteriori inference of these performances. Both highlight the same at-risk dynamic regime. Therefore, our UQ framework is useful for fine-tuning, adaptive training, or error detection.
@@ 4.4 Robustness across the parametric regime
The Continuous Ranked Probability Score (CRPS) is on the order of $10^{-4}$, which is extremely low. Conceptually, CRPS measures the difference between the mean distance of the ensemble from the true target and the variance within the ensemble. A low CRPS indicates that each sampled trajectory is close to the true target trajectory and that the ensemble members themselves are close to each other. Since the model performs well near the training distribution, there is little variation in the ensemble predictions. However, as performance decreases (e.g., when $R_{e}>130$), the CRPS increases, reflecting a greater variety within the ensemble. Such a low CRPS is synonymous with an accurate and reliable model.
@@
The results show that expanding the training distribution reduces the relative MSE to below 0.2% on test data for the entire parameter range of interest, which is, to our knowledge, is among the lowest relative MSE reported for the similar benchmarks of the “flow around a cylinder”. The CRPS remains low, but has not decreased as drastically as the relative MSE, indicating that while the model has become more accurate, the diversity of the ensemble has increased. As the dynamic range in the training distribution expands, the model becomes aware of a wider range of possible dynamics, leading to a decrease in overall confidence. However, in section 5 we show that the aggregated uncertainty is not as relevant as its distribution over the parameter space.
5 Adaptive Sampling for Optimized Generalization
@@
A parametrised model naturally prompts the question of how to select the appropriate parameters to ensure effective learning and inference across the entire parameter space. Since gathering training data for the full domain is often infeasible, we are typically limited to a small subset of the parameter space. The central issue, therefore, is identifying the minimal set of parameters that will offer the most comprehensive information, enabling the model to generalize effectively across the entire range of parameters. We aim to use Uncertainty Quantification (UQ) to efficiently explore the parameter space and construct an adaptive sampling strategy. To this end, the high uncertainties identify regions where model performance is suboptimal and indicate the region of the parameter space where data needs to be collected. This approach leads to an adaptive sampling technique that adjusts data sampling based on model performance or uncertainty to improve learning efficiency and accuracy. Our approach is implemented during inference of the UQ framework and reduces the reliance on prior knowledge of the parameter space and the need for validation data throughout the adaptive procedure, which would be necessary with performance/error driven adaptive sampling.
@@
This approach can only be effective if the uncertainty can serve as a reliable measure of model performance or predictability. To validate this, we first examine whether there is a strong correlation between the a posteriori error and the a priori uncertainty. To do this, we choose a parametrised model with a variable Reynold number $\xi=Re$ and a fixed aspect ratio $(\beta=1.35)$. This configuration leads to a bifurcation at Reynolds number $\approx 60$. The dynamical regime before this critical Reynolds number $(R_{ec})$ shows a stationary behaviour, while the regime after the bifurcation results in a time-varying limit cycle, as illustrated in Figure 3 and in Figure 13 by the different colours of the shaded regions. The experiment is performed over a parameter range from $R_{e}=50$ in the pre-bifurcation regime to $R_{e}=140$ in the post-bifurcation regime. The model infers over this entire range on 10 equally spaced Reynolds number values, where we have validation data. We obtain 10 measurements of $\nu_{\xi}$ which we compare with the model’s performance over the same range. Performance is evaluated using the Mean Squared Error (MSE), scaled by the order of magnitude of the signal over the range. Note that in real-world scenarios, validation data over the full range would not be necessary. The scaled MSE is computed here a posteriori solely to assess and validate our uncertainty quantification.
@@
At initialisation, the model is trained with only two samples in the post-bifurcation regime: $R_{e}=90$ and $R_{e}=120$, exposing the model to minimal parameter variety. Figure 13-a shows the initial uncertainty distribution over the parameter range $\nu_{\xi}$, while Figure 13-b shows the scaled error. There is a strong correlation between the two values as formally shown in 13-c, with the linear fit between $\nu_{\xi}$ and the error, along with the corresponding Pearson correlation index. In this first step, we observe a Pearson index of $r=0.87>0.5$, indicating a strong, positive linear correlation. Not surprisingly, the model is most confident for the in-distribution parameters and specifically for the two training samples. It is least confident in the pre-bifurcation regime. Since the model has never encountered this regime, it has both higher uncertainty and lower performance. This analysis confirms that the UQ measure serves as a reliable indicator of model predictability or error, and provides the basis for an adaptive sampling strategy that iteratively guides the retraining of the model until uncertainty is minimised over the parameter range of interest.
@@
The uncertainty distribution given by $\nu_{\xi}$ in Figure 13-a shows the highest uncertainty at $R_{e}=50$. Therefore, we feed the model with this data and perform a first retraining. This retraining is significantly shorter than the initial training, with fewer initial samples included to avoid forgetting. After retraining, both uncertainty and MSE are recomputed at inference and a posteriori, respectively, and show a complete redistribution across the parameter range, as shown in Figure 13-d,e.
@@
After the first retraining, the confidence in the pre-bifurcation segment improves significantly. The stationary-like dynamics make inference easier for the model as the output does not change with time. At this stage, error and uncertainty remain well correlated, as shown in Figure 13-f. The highest uncertainty now shifts to the other side of the bifurcation, at $R_{e}=80$, suggesting the need to refine the model’s understanding of the segment near the critical Reynolds number. The data associated with this parameter value becomes the new model’s proposal for retraining. Interestingly, at this stage, extrapolation beyond the training distribution ($R_{e}>120$) is less challenging in terms of both uncertainty and error compared to the lower range ($R_{e}<90$). The retraining process can be extended by iteratively adding the samples with the highest uncertainty until the uncertainty converges below a desired value. The complete iterative retraining process, together with the uncertainty distribution $\nu_{\xi}$, the scaled MSE, and their correlations is shown in Figure 13. The results throughout this process indicate that points with the highest Mean Squared Error (MSE) consistently align with those exhibiting the highest uncertainty. A clear and consistent linear relationship between error and uncertainty is evident. Following each retraining iteration, the Pearson correlation coefficient between the error and MSE exceeds 0.7, underscoring a strong positive linear correlation. This demonstrates the effectiveness of our uncertainty quantification method in anticipating poor performance across the parameter space and for adaptive sampling.
@@
After completing five retraining iterations, the UP-dROM method demonstrates consistent and uniform learning performance across the entire range of model parameters, effectively capturing the underlying bifurcation dynamics. Notably, the model’s uncertainty appears to have converged to a stable final distribution, denoted as $\nu_{\xi}$. The model progressively identifies parameter points that minimize the error at each training step. As retraining continues at these optimal points, we observe a convergence of both the scaled mean squared error (MSE) and the uncertainty, which both tend toward an optimal distribution.
6 Conclusion

### v3

@@ 4 Uncertainty quantification
Most conventional reduction strategies rely on deterministic estimation, where the latent space projection is represented by a single point estimate. This approach fails to capture the inherent uncertainty within the model. In contrast, UP-dROM utilises variational methods to assess uncertainties, providing a probabilistic projection in the latent space instead of a deterministic one. This is achieved through the use of a Variational Autoencoder (VAE). Specifically, given the observed data $X=D$ (i.e., the dataset $D$), the method seeks to infer the posterior distribution $p(z|X=D)$, where $z$ represents the latent variables.
The central idea connecting variational inference and uncertainty quantification revolves around approximating the true posterior distribution $p(z|X=D)$ using a surrogate distribution $q(z|X=D)$. In this case, the surrogate is chosen to be a Gaussian distribution due to its desirable mathematical properties and ease of interpretation. To ensure that the surrogate closely approximates the true posterior, we minimise the Kullback-Leibler Divergence (KLD) between $p(z|X=D)$ and $q(z|X=D)\sim\mathcal{N}(\mu,\sigma^{2})$, as:
@@
Since $p(z|X=D)$ is not known a priori, we instead utilise the Evidence Lower Bound (ELBO), which can be computed at any stage. Variational inference theory states that minimizing the ELBO, also minimises the KLD between the true posterior and the surrogate distribution [31]. The ELBO is given by:
@@
where $p(z)$ is the prior Gaussian distribution representing the distribution of the latent space. It is typically chosen as $p(z)\sim\mathcal{N}(0,I)$, which imposes a standard normal prior on the latent variables. This choice ensures that the latent space is continuous and well-structured. It also simplifies the mathematical treatment of the model, making the optimisation more tractable. The KLD between two Gaussian distributions such as $q(z|X=D)$ and $p(z)$ can be computed as follows,
@@ 4.1 Uncertainty in space
We can use the uncertainty measure $\nu_{d,t,\xi}$ to visualise how uncertainty is distributed across the computational domain at different time steps for a fixed parametrisation. Figure 9-a shows the $u$-velocity fields at three key time steps: (1) before the wake forms, during the unstable stationary-like behaviour, (2) during the transition, and (3) once the system has reached the stable limit-cycle, as highlighted in the kinetic energy plot. Figure 9-b illustrates the distribution of uncertainty in space $\nu_{d}$ at these same time steps, computed as:
@@
Figure 9: U velocity field at the selected snapshots (a), the corresponding uncertainty fields $\nu_{d}$ (b), and the snapshot locations indicated on the kinetic energy signal (c), for a fixed parametrisation $\xi$: $R_{e}=90$.
This figure illustrates a strong correlation between the known space-time dynamics and the uncertainty measurement. Prior to the system’s transitioning to the stable limit cycle, uncertainty is evenly distributed along the symmetry axis of the flow domain, as expected, but this distribution breaks down as the dynamics transition. It is in this transient regime that the uncertainty across the entire domain is the highest. The highest confidence is observed upstream of the obstacle, where the dynamics remain relatively stationary throughout the time window. In contrast, lower confidence is concentrated at the edges of the obstacle’s wake, suggesting that uncertainty is more closely related to fluctuations in the flow, rather than the instantaneous flow values in specific regions. These fluctuations increase variability within the system, leading to higher uncertainty.
@@ 4.2 Uncertainty in Time
Notably, the true signal lies almost entirely within the confidence interval. It is important to emphasise that this Uncertainty Quantification (UQ) strategy is not time-dependent and is solely driven by the input state. If it were time-dependant, the uncertainty would gradually increase as the dynamic process evolves autoregressively. Instead, we propose a likelihood measure for obtaining a given value, independent of previous values. This approach provides an instantaneous uncertainty measure with minimal computational cost.
As shown in Figure 10, the greatest uncertainties occur at the peaks of the signal, corresponding to instances when the local dynamics are most critical. This typically happens when the obstacle’s wake shifts within the flow domain, causing substantial changes that the model must predict with accuracy. These transitions trigger increased uncertainty. This behaviour has a significant physical interpretation: uncertainty is highest in regions and time periods where the flow dynamics are most sensitive to changes, such as during sharp transitions, shifts in wake structure, small-scale patterns, and rare events.
@@
Furthermore, as expected, higher uncertainty is observed in the transient part of the signal, compared to the limit cycle. This phenomenon is formally illustrated in Figure 11. To analyse temporal variations in uncertainty for a fixed parameter, we compute the spatially averaged uncertainty $\nu_{t}$ at each time step $t$ over $n_{xy}$ spatial locations,
@@
The width of the bell curve in Figure 11 aligns with the transient phase of the dynamic evolution visible on the kinetic energy signal. Starting from the initial condition, the uncertainty increases as the system transitions towards the limit cycle. Once this stable state is reached, the uncertainty decreases and stabilises at a plateau. This can be explained by two factors: (1) the regime transition induces variability, which like any other source of variability, is associated with lower confidence, and (2) the training distribution contains fewer transient states compared to the limit cycle. As a result, the model has been trained more extensively on the stable state than on the transient phases.
@@ 4.3 Uncertainty in the Parameter space
Uncertainty Quantification also provides a confidence measure across the parameter space, aggregating the spatial and temporal components,
@@
with $n_{t}$ the time discretisation and $\bar{\phi}_{\xi,d,t}$ the mean of the ensemble of size $n$.
As shown in Figure 7, the model can learn and infer within the 2D parameter space. Specifically, it can interpolate within the training subspace and extrapolate with varying accuracy depending on the region of the parameter space. The further the inference distribution deviates from the training distribution, the greater the uncertainty of the model, although not uniformly. The direction in which the inference distribution moves away from the training distribution affects the performance of the model. Our uncertainty quantification method can measure this non-Euclidean distance. Figure 12 illustrates how our uncertainty quantification process allows us to identify these regions during inference, thus reducing the need for simulation data for evaluation. To improve the performance of the model, it could be retrained using samples from and only from these uncertain regions. An important observation is the strong agreement between the a-posteriori performance evaluation shown in Figure 7 and the a-posteriori inference of these performances. Both highlight the same at-risk dynamic regime. Therefore, our UQ framework is useful for fine-tuning, adaptive training, or error detection.
@@ 4.4 Robustness across the parametric regime
The Continuous Ranked Probability Score (CRPS) is on the order of $10^{-4}$, indicating a very low value. Conceptually, CRPS measures the difference between the mean distance of the ensemble from the true target and the variance within the ensemble. A low CRPS indicates that each sampled trajectory is close to the true target trajectory and that the ensemble members themselves are close to each other. Since the model performs well near the training distribution, there is little variation in the ensemble predictions. However, as performance decreases (e.g., when $R_{e}>130$), the CRPS increases, reflecting a greater variety within the ensemble. Such a low CRPS is synonymous with an accurate and reliable model.
@@
The results show that expanding the training distribution reduces the relative MSE to below 0.2% on test data for the entire parameter range of interest, which, to our knowledge, is among the lowest relative MSE reported for the similar benchmarks of the “flow around a cylinder”. The CRPS remains low, but has not decreased as drastically as the relative MSE, indicating that while the model has become more accurate, the diversity of the ensemble has increased. As the dynamic range in the training distribution expands, the model becomes aware of a wider range of possible dynamics, leading to a decrease in overall confidence. However, in Section 5 we show that the aggregated uncertainty is not as relevant as its distribution over the parameter space.
5 Adaptive Sampling for Optimised Generalisation
@@
A parametrised model naturally prompts the question of how to select the appropriate parameters to ensure effective learning and inference across the entire parameter space. Since gathering training data for the full domain is often infeasible, we are typically limited to a small subset of the parameter space. The central issue, therefore, is identifying the minimal set of parameters that will offer the most comprehensive information, enabling the model to generalise effectively across the entire range of parameters. We aim to use Uncertainty Quantification (UQ) to efficiently explore the parameter space and construct an adaptive sampling strategy. To this end, the high uncertainties identify regions where model performance is suboptimal and indicate the region of the parameter space where data need to be collected. This approach leads to an adaptive sampling technique that adjusts data sampling based on model performance or uncertainty to improve learning efficiency and accuracy. Our approach is implemented during inference of the UQ framework and reduces the reliance on prior knowledge of the parameter space and on the need for validation data throughout the adaptive procedure, which would be necessary with performance/error driven adaptive sampling.
@@
This approach can only be effective if the uncertainty can serve as a reliable measure of model performance or predictability. To validate this, we first examine whether there is a strong correlation between the a posteriori error and the a priori uncertainty. To do this, we choose a parametrised model with a variable Reynolds number $\xi=Re$ and a fixed aspect ratio $(\beta=1.35)$. This configuration leads to a bifurcation at Reynolds number $\approx 60$. The dynamical regime before this critical Reynolds number $(R_{ec})$ shows a stationary behaviour, while the regime after the bifurcation results in a time-varying limit cycle, as illustrated in Figure 3 and in Figure 13 by the different colours of the shaded regions. The experiment is performed over a parameter range from $R_{e}=50$ in the pre-bifurcation regime to $R_{e}=140$ in the post-bifurcation regime. The model infers over this entire range on 10 equally spaced Reynolds number values, for which validation data are available. We obtain 10 measurements of $\nu_{\xi}$ that we compare with the model’s performance over the same range. Performance is evaluated using the Mean Squared Error (MSE), scaled by the order of magnitude of the signal over the range. Note that in real-world scenarios, validation data over the full range would not be necessary. The scaled MSE is computed here a posteriori solely to assess and validate our uncertainty quantification.
@@
At initialisation, the model is trained with only two samples in the post-bifurcation regime: $R_{e}=90$ and $R_{e}=120$, exposing the model to minimal parameter variety. Figure 13-a shows the initial uncertainty distribution over the parameter range $\nu_{\xi}$, while Figure 13-b shows the scaled error. There is a strong correlation between the two values as formally shown in Figure 13-c, with the linear fit between $\nu_{\xi}$ and the error, along with the corresponding Pearson correlation index. In this first step, we observe a Pearson index of $r=0.87>0.5$, indicating a strong, positive linear correlation. Not surprisingly, the model is most confident for the in-distribution parameters and specifically for the two training samples. It is least confident in the pre-bifurcation regime. Since the model has never encountered this regime, it exhibits both greater uncertainty and poorer performance. This analysis confirms that the UQ measure serves as a reliable indicator of model predictability or error, and provides the basis for an adaptive sampling strategy that iteratively guides the retraining of the model until uncertainty is minimised over the parameter range of interest.
@@
The uncertainty distribution given by $\nu_{\xi}$ in Figure 13-a shows the highest uncertainty at $R_{e}=50$. Therefore, we feed the model with this data and perform the first retraining. This retraining is significantly shorter than the initial training, with fewer initial samples included to avoid forgetting. After retraining, both uncertainty and MSE are recomputed at inference and a posteriori, respectively, and exhibit a complete redistribution across the parameter range, as shown in Figure 13-d,e.
@@
After the first retraining, the confidence in the pre-bifurcation segment improves significantly. The stationary-like dynamics make inference easier for the model as the output does not change with time. At this stage, error and uncertainty remain well correlated, as shown in Figure 13-f. The highest uncertainty now shifts to the other side of the bifurcation, at $R_{e}=80$, suggesting the need to refine the model’s understanding of the segment near the critical Reynolds number. The data associated with this parameter value becomes the next candidate point for retraining. Interestingly, at this stage, extrapolation beyond the training distribution ($R_{e}>120$) is less challenging in terms of both uncertainty and error compared to the lower range ($R_{e}<90$). The retraining process can be extended by iteratively adding samples with the highest uncertainty until the uncertainty converges below a desired value. The complete iterative retraining process, together with the uncertainty distribution $\nu_{\xi}$, the scaled MSE and their correlations is shown in Figure 13. The results throughout this process indicate that points with the highest Mean Squared Error (MSE) consistently align with those exhibiting the highest uncertainty. A clear and consistent linear relationship between error and uncertainty is evident. Following each retraining iteration, the Pearson correlation coefficient between the UQ and MSE exceeds 0.7, underscoring a strong positive linear correlation. This demonstrates the effectiveness of our uncertainty quantification method in anticipating poor performance across the parameter space and for adaptive sampling.
@@
After completing five retraining iterations, the UP-dROM method demonstrates consistent and uniform learning performance across the entire range of model parameters, effectively capturing the underlying bifurcation dynamics. Notably, the model’s uncertainty appears to have converged to a stable final distribution, denoted as $\nu_{\xi}$. The model progressively identifies parameter points that minimise the error at each training step. As retraining continues at these optimal points, we observe a convergence of both the scaled mean squared error (MSE) and the uncertainty, which both tend towards an optimal distribution.
6 Conclusions and perspectives